OpenStudy (pinkkjazzy3):

Write an expression in (a+bi)+(c+di) form involving the addition of two complex numbers, where a, b, c, and d are all positive integers and where the sum is 4+3i.

OpenStudy (pinkkjazzy3):

plzzzz!!!

OpenStudy (anonymous):

\[(a+bi)+(c+di)=(a+c)+(b+d)i\] so if \[(a+bi)+(c+di)=4+3i\] that tells you \[a=c=4,b+d=3\]

OpenStudy (anonymous):

oops typo there, i meant \[a+c=4,b+d=3\]